You have given a function λ:R-> R with the following properties (x ∈ R, n∈ N) :
Λ(n) =0, λ(x+1)= λ(x), λ(n+1/2)=1
Find two functions p,q:R-> Rwith q(x) not equal to 0 for all x such that λ(x)= q(x)(p(x)+1)
1
Expert's answer
2020-11-17T07:35:23-0500

Let "x" be a real number.The floor function "\lfloor x \rfloor" is defined to be the greatest integer less than or equal to the real number "x". The fractional part function "\{ x \}"  is defined to be "\{x\}= x -\lfloor x \rfloor". Define two functions "p,q:\mathbb R\to\mathbb R" with "q(x)" not equal to 0 for all "x" in the following way: "p(x)=2\{x\}-1, q(x)=\begin{cases}4\{x\}-1,\ x\ne\frac{1}{4}+n,n\in\mathbb N \\2,\ x=\frac{1}{4}+n,n\in\mathbb N\end{cases}".


Then the function "λ:\mathbb R\to\mathbb R," "λ(x)= q(x)(p(x)+1)=\begin{cases}(4\{x\}-1)(2\{x\}), \ x\ne\frac{1}{4}+n,n\in\mathbb N\\ 1,\ \ \ x=\frac{1}{4}+n,n\in\mathbb N\end{cases}" has the following properties:


"\lambda(n)=(4\{n\}-1)(2\{n\}) =-1\cdot0=0,\ \ \lambda(x+1)= \lambda(x)," and


"\lambda(n+\frac{1}{2})=(4\{n+\frac{1}{2}\}-1)(2\{n+\frac{1}{2}\})=(4\cdot\frac{1}{2}-1)(2\cdot\frac{1}{2})=1"


for all "x \in\mathbb R, n\in\mathbb N."